Using the force method, Determine the reactions at supports. EI is constant.

Determine the reactions at supports. El is constant. 500 lb/ft 3k 10 ft- SA 0 a 6.5 (10) 10 ft - 2.5 uz
Answer #1

PACE NO.: DATE: 1 Solution : - b - a AL oskift loft - MA loft use I lb = 0.001k heb Apply minimum potential (strain energy mePAGE !0.: DATE: 1 Mr = Cy (x) - 0.5 (x) (ч.) o3 x < 10 UBA = CM² 2. І єг — JUBA 2LMх а = . 2ЕГ DUBA Әсч 2 vед - Гусе) – без гPAGE NO.: | DATE: 1 1 - Applying equilibrium equation *_F@Dof_BA_spam, weget_4e- By = 1*29Sk MA Hх *= 0 , P tex =e, - (2 м),